OpenStudy (anonymous):

What is the arc measure of a sector with area 38.485 ft2 and radius 7?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

First find the area of the entire circle

jimthompson5910 (jim_thompson5910):

A = pi*r^2

jimthompson5910 (jim_thompson5910):

Then you use the formula P = (t/360)*A where... P = area of the piece of the circle (or the sector area) t = central angle (also the arc measure) A = area of the entire circle

OpenStudy (anonymous):

thanks jim :)

jimthompson5910 (jim_thompson5910):

you're welcome